FY16 Factory Rent $330,000 Direct Materials Inventory, Beginning 96,000 Direct Materials Inventory, Ending 87,000 Direct...

   FY16
Factory Rent $330,000
Direct Materials Inventory, Beginning 96,000
Direct Materials Inventory, Ending 87,000
Direct Materials Purchases 654,000
Direct Labor-Wages 425,000
Indirect Labor-Wages 28,000
Finished Goods Inventory, Beginning 25,000
Finished Goods Inventory, Ending 44,000
Indirect Materials 66,000
Plant Utilities 40,000
General and Administrative 101,350
Work-in-Process Inventory, Beginning 27,000
Work-in-Process Inventory, Ending 33,000
Marketing Expenses 225,000
Sales Revenue 2,550,000
A.   Prepare a statement of cost of goods manufactured and an income statement for the year ended December 31, 2016.
B. For FY2016, calculate gross margin %. What is Jamison Tools largest expense as a % of revenue?
C. Interpret the results of the financial statements
Answer #1

A.

                                       Jamison Tools

              Statement of Cost of Goods Manufactured

            For the for the year ended December 31,2016

Opening Inventory of Direct Material

96000

Add: Purchases of Direct Materials

654000

Less: Closing Inventory of Direct Material

(87000)

663000

Direct Labour-Wages

425000

Prime Cost

1,088,000

Add: Factory Overhead Cost

4,64,000

Total Factory Cost

1,552,000

Add: Opening work-in-process

   27,000

Less:Closing work-in-process

   (33000)

Cost of Goods Manufactured

1,546,000

Working note:Calculation of factory overheads:

Factory rent       $ 330000

Indirect labour $   28000

Indirect material $66000

Plant utilities      $ 40000

TOTAL $464000

                      Jamison Tools

Income statement for the year ended December 31,2016

Sales Revenue

2,550,000

Cost of Goods Sold

Cost of Goods Manufactured

1,546,000

Add: Opening Finished Goods Inventory

      25,000

Goods Available for sale

1,571,000

Less: Closing Finished Goods Inventory

       44,000

Gross Margin

1,527,000

Less: Administrative Overheads

    (101350)

Less: Marketing Expenses

   (225000)

Income Before Taxes

1,200,650

B. Gross Margin Percentage=(1,527,000*100)/ 2,550,000=59.8823%

    Largest Expense as a percentage of Revenue=Expense of Direct Material

                                                                  =654000*100/2550000=25.647 %

Calculation of factory overheads:

Factory rent       $ 330000

Indirect labour $   28000

Indirect material $66000

Plant utilities      $ 40000

TOTAL                   $464000